Question 5: Suppose market demand for poutine is Q=75−P.

(a) If Claude stops making political donations, how many poutineries do we expect to operate in long run competitive equilibrium? (Assume that the cost function from Q4 remains at C(q) =3q+4q2+64).

(b) What will the market price be?

(c) How much total poutine will be produced?

(d) What will consumer surplus be?
Answer #1

Marginal cost (MC) = dC(q)/dq = 3 + 8q

Average cost (AC) = C(q)/q = 3 + 4q + (64/q)

(a) In long run equilibrium, P = MC = AC.

3 + 8q = 3 + 4q + (64/q)

4q = 64/q

q2 = 64/4 = 16

q = 4

P = MC = 3 + (8 x 4) = 3 + 32 = 35

From demand function, Q = 75 - 35 = 40 (Market quantity produced)

Number of firms = Q/q = 40/4 = 10

(b) Market price = 35

(c) Total (market) quantity (Q) = 40

(d) From demand function, when Q = 0, P = 75 (Vertical intercept)

Consumer surplus = Area between demand curve & price = (1/2) x (75 - 35) x 40 = 20 x 40 = 800